OpenStudy (anonymous):

Which is equal to (x + 4)(3x2 4x + 2)? (Points : 1) 3x^3 - 16x^2 - 14x + 8 3x^2 - 4x + 8 3x^2 3x + 6 3x^3 4x^2 + 2x + 8

OpenStudy (anonymous):

x (3x^2 - 4x + 2) + 4 = (3x^3 - 4x^2 + 2x) + 4 4 (3x^2 - 4x + 2) = (12x^2 - 16x + 8)

OpenStudy (anonymous):

together: (3x^3 - 4x^2 + 2x) + (12x^2 - 16x + 8) 3x^3 - 4x^2 + 12x^2 + 2x - 16x + 8 3x^3 - 16x^2 -14x + 8 (this should be the answer)

OpenStudy (anonymous):

it was just to much to put down to simplify it by grouping lol

OpenStudy (anonymous):

otherwise it would look like this: (+3x^3) (-) (- 4x^2) (+) (+12x^2) (+) (+2x) (-) (-16x) (+) (+8) that is what my teacher did not show us

OpenStudy (anonymous):

\[(x+4)(3x^{2}-4x + 2)\] \[= x(3x^{2} -4x + 2) + 4(3x^{2} -4x + 2)\] \[= (3x^{3} - 4x^{2} + 2x) + (12x^{2} - 16x + 8)\] Redistributing. \[= 3x^{3} + 8 x^{2} - 14 x + 8\]

OpenStudy (anonymous):

i dont have that choice though

OpenStudy (anonymous):

That's odd. It's the correct answer. :/

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=+%28x+%2B+4%29%283x2+-4x+%2B+2%29 Check the first alternate form.

OpenStudy (anonymous):

i know how they groped it, they did it the same way you did but how i did mine idont know either

OpenStudy (anonymous):

this is algebra2 i think

OpenStudy (anonymous):

are they in order from greatest to least?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

oh it is : (x + 4)(-3x2 - 4x + 2) so: -3x^3 - oh dang this is all wrong

OpenStudy (anonymous):

(x + 4)(-3x2 - 4x + 2) 1. x ( -3x^2 - 4x + 2) + 4 ( -3x^2 - 4x + 2) 2. (-3x^3 -4x^2 + 2x) + (-12x^2 - 16x + 8) 3. -3x^3 - 4x^2 -12x^2 + 2x - 16x + 8 4. -3x^3 - 16x^2 -14x + 8

OpenStudy (anonymous):

its correct